# Global minimality of the Hopf map in the Faddeev–Skyrme model with large coupling constant

### v1

@@ 2.3. The Hopf map and horizontal conformality
In this subsection we gather some well-known facts about the Hopf map. We first recall the following definition, which the reader can find more generally in [3, Definition 2.4.2 and Lemma 2.4.4]:
@@
$du(x)\circ du(x)^{t}=\frac{1}{2}|du(x)|^{2}\,\mathrm{Id}_{T_{u(x)}{\mathbb{S}}^{2}}\quad\text{for }\mathcal{H}^{3}\text{-a.e. }x\in{\mathbb{S}}^{2}\,,$ (2.13)
@@
The following lemma gives a simple but very useful pointwise inequality related to weak horizontal conformality. Its proof is elementary and can be found for instance in [17, Proof of Lemma 2.1].
@@
We remark that this type of estimate also plays an important role in the minimality arguments in [7].
@@
One can then choose an orthonormal frame $\{\tau_{1},\tau_{2},\tau_{3}\}$, defined at a point $(t,e^{i\varphi_{1}},e^{i\varphi_{2}})\in{\mathbb{S}}^{3}$ as
@@ 3. The Hopf invariant for maps of finite Faddeev–Skyrme energy
In this section we prove some results concerning the definition and integrality of the Hopf invariant for maps with finite Faddeev–Skyrme energy, cf. (1.2). The integrality property of the Hopf invariant in this class was correctly claimed in [22, Section 3] but, as explained in [2, §2.1.4], there is a gap in the argument. In [2, Corollary 5.2] a complete proof of integrality was given, but the proof therein is fairly long and not self-contained, as it relies on the earlier results by the authors in [1]. Here we present a new, more concise proof of this integrality which, although also not self-contained, we believe to be fairly transparent, and which contains ingredients which will also be important in the sequel.
@@
and in view of the fact that $H^{2}_{\rm{dR}}({\mathbb{S}}^{3})=0$ there exists $\beta\in\Omega^{1}({\mathbb{S}}^{3})$ such that $u^{*}\omega_{{\mathbb{S}}^{2}}=d\beta$. The Hopf invariant or Hopf charge of $u$, see [34], is then given by the integral formula (1.3), and in particular $Q(u)\in\mathbb{Z}$.
@@
$\mathcal{A}_{3}({\mathbb{S}}^{3}):=\big{\{}v\in W^{1,2}({\mathbb{S}}^{3};{\mathbb{S}}^{3})\colon\textup{cof}(dv)\in L^{3/2}({\mathbb{S}}^{3})\big{\}}\,,$ (3.1)
where $\textup{cof}(dv)$ is identified with the pull-back map $v^{*}\colon\Omega^{2}({\mathbb{S}}^{3})\to\Omega^{2}({\mathbb{S}}^{3})$. These function spaces were introduced in the context of nonlinear elasticity theory in [4] and appear naturally in our problem, since we have the following generalization of the results in [10, 26, 32], due to [13, Theorems 1.3 and 4.4]:
@@
$Q(u)=\frac{1}{16\pi^{2}}\int_{\mathbb{S}^{3}}\beta\wedge d\beta=\deg(\hat{u})\in\mathbb{Z}\,.$ (3.4)
In addition to Proposition 3.2, the following result due to [5] is the key ingredient in our proof of Theorem 3.3.
@@
By [5, Theorem 1], a map $u\colon{\mathbb{S}}^{3}\to{\mathbb{S}}^{2}$ is $W^{1,2}$-strongly approximable by smooth maps if and only if it pulls-back (distributionally) closed $2$-forms on ${\mathbb{S}}^{2}$ to (distributionally) closed 2-forms on ${\mathbb{S}}^{3}$. Since any form in $\Omega^{2}(\mathbb{S}^{2})$ can be represented as $f\omega_{\mathbb{S}^{2}}$ for some $f\in C^{\infty}({\mathbb{S}}^{2})$, it is then necessary and sufficient to check closedness of $u^{*}\omega_{\mathbb{S}^{2}}$. ∎
@@
This is detailed in [2, Lemma 3.2]. Fix $\{j,k,\ell\}\subset\{1,2,3\}$. Thanks to classical approximation by (localisation and) mollification, we see that $u^{k},u^{\ell}\in W^{1,2}({\mathbb{S}}^{3};\mathbb{R})$ satisfy
@@
hence the 2-form $\alpha^{k\ell}\colon=du^{k}\wedge du^{\ell}$ is distributionally closed, and using again approximation by mollification of $u^{j}\in W^{1,2}({\mathbb{S}}^{3};\mathbb{R})$ and of the closed form $\alpha^{k\ell}\in L^{2}({\mathbb{S}}^{3};\bigwedge^{2}T^{*}{\mathbb{S}}^{3})$, we infer that
@@
$d(u^{*}\omega_{\mathbb{S}^{2}})=3\,du^{1}\wedge du^{2}\wedge du^{3}\in L^{1}\Big{(}{\mathbb{S}}^{3};\bigwedge\nolimits^{3}T^{*}{\mathbb{S}}^{3}\Big{)}\,,$
@@
Since $u(x)\in\mathbb{S}^{2}$ for $\mathcal{H}^{3}$-a.e. $x\in{\mathbb{S}}^{3}$, the vectors $(du^{i}(x))_{i=1}^{3}$ must be linearly dependent, hence the last wedge product vanishes. ∎
The next lemma, due to [14], gives us a gauge for the lifts of $u$.
@@
$h^{*}\omega_{{\mathbb{S}}^{2}}=2d\theta\,.$ (3.6)
@@
The proof in the case of a smooth map $u$ can be found in [14, Lemma 2.1], and otherwise the conclusion of (3.7) can be repeated analogously, since the calculations therein can be carried over pointwise $\mathcal{H}^{3}$-a.e. on ${\mathbb{S}}^{3}$. The coefficients in (3.7) are actually different in [14] due to a different choice of normalization for the metric on $\mathbb{S}^{2}$ (which also changes the normalization constant in the integral formula (1.3) for the Hopf degree). For the readers’ convenience we reproduce the proof of (3.7) with the metric used here. In the orthonormal frame $\{\tau_{1},\tau_{2},\tau_{3}\}$ on $T{\mathbb{S}}^{3}$, the $1$-form $\theta$ corresponds to the scalar product against $\tau_{1}$. Thus, the condition $\beta=2\hat{u}^{*}\theta$ and (2.18)–(2.19), imply
@@
$u_{j}\rightarrow u\text{ strongly in }W^{1,2}({\mathbb{S}}^{3};{\mathbb{S}}^{2})\text{ as }j\to\infty\,.$ (3.10)
@@
since $\frac{5}{4}<\frac{3}{2}=:1^{*}$. We note that the borderline $L^{3/2}$-estimate is actually also true, see e.g. [20] or [6, Corollary 12], at least in $\mathbb{R}^{3}$ instead of $\mathbb{S}^{3}$. Invoking Lemma 3.6, we now fix lifts $(\hat{u}_{j})_{j\in\mathbb{N}}\subset C^{\infty}({\mathbb{S}}^{3};{\mathbb{S}}^{3})$ satisfying
@@
cf. (3.5). In view of (2.14) and (3.10), by passing to a non-relabelled subsequence, we obtain $\beta\in L^{5/4}\big{(}{\mathbb{S}}^{3};\bigwedge\nolimits^{1}T^{*}{\mathbb{S}}^{3}\big{)}$ so that
@@
Indeed, in view of (2.15), the forms $\hat{u}_{j}^{*}\theta$ can be written in coordinates as second-order polynomials of the form $\hat{u}_{j}\bullet d\hat{u}_{j}$, where the latter notation indicates that each monomial in the expression is of first order separately in $\hat{u}_{j}$ and $d\hat{u}_{j}$. In particular, since $\hat{u}_{j}\rightarrow\hat{u}$ strongly in $L^{p}({\mathbb{S}}^{3};{\mathbb{S}}^{2})$ for every $1\leq p\leq 6$ and $d\hat{u}_{j}\rightharpoonup d\hat{u}$ weakly in $L^{5/4}({\mathbb{S}}^{3};\bigwedge\nolimits^{1}T^{*}{\mathbb{S}}^{3})$, this particular product structure of weakly convergent objects and strongly convergent ones justifies (3.17).
@@
where $\{f_{2},f_{2}\}$ is the orthonormal frame on $T{\mathbb{S}}^{2}$ defined after (2.20). Thus, using also (2.4), we have
@@
Since $\hat{u}\in\mathcal{A}_{3}({\mathbb{S}}^{3})$, we can apply Proposition 3.2 to infer that $\deg(\hat{u})$ is a well-defined integer and that $\hat{u}^{*}$ and $d$ commute. Hence, recalling (2.16), we have
@@
$=\frac{1}{4\pi^{2}}\int_{\mathbb{S}^{3}}\hat{u}^{*}\theta\wedge\hat{u}^{*}(d\theta)=\frac{1}{4\pi^{2}}\int_{\mathbb{S}^{3}}\hat{u}^{*}(\theta\wedge d\theta)=\frac{1}{2\pi^{2}}\int_{\mathbb{S}^{3}}\hat{u}^{*}\omega_{{\mathbb{S}}^{3}}=\deg(\hat{u})\,,$
@@
There is a serious obstruction to having a more direct proof of Theorem 3.3, namely that it is not known whether one can approximate in the Faddeev–Skyrme energy a general map in $\mathcal{U}_{\mathcal{FS}}$ by smooth maps. This is not known even if one ignores the image constraint on maps in $\mathcal{U}_{\mathcal{FS}}$, and we refer the reader to [12] for some partial results in the unconstrained case. We emphasize that from the above proof one cannot infer that the Hopf invariant of $u$ is obtained as the limit of the Hopf invariants of $(u_{j})_{j\in\mathbb{N}}$, as in (3.10), since the latter sequence converges to $u$ only in too weak a sense for this purpose (and similarly for the liftings).
@@
Since the map $H^{2}_{\mathrm{dR}}({\mathbb{S}}^{2})\ni[\omega]\mapsto\int_{{\mathbb{S}}^{2}}\omega\in\mathbb{R}$ is a linear isomorphism, our assumption in particular implies that
$[\omega_{1}]=[\omega_{2}]\in H_{\mathrm{dR}}^{2}({\mathbb{S}}^{2})\,,$
@@
[…]
@@
$\int_{{\mathbb{S}}^{2}}\omega_{1}=\int_{{\mathbb{S}}^{2}}\omega_{2}\,.$ (3.26)
@@ 4. Relaxed form of the perturbed Faddeev–Skyrme energy
In order to prove Theorem 1.1, and following the approach in [17, 27], we consider a relaxation of $\mathcal{FS}_{\rho}$ to the space of closed differential forms. Let us first define, for a closed 2-form $\alpha$ on ${\mathbb{S}}^{3}$,
@@
By Stokes’ Theorem this quantity, referred to as the Hopf invariant of the form $\alpha$, is well-defined, i.e.$Q$ depends on $\beta$ only through $d\beta$, see also the corresponding discussion in Remark 3.1 for the Hopf invariant of a map $u$. Introducing the space of admissible forms
$\mathcal{A}_{\mathcal{FS}}:=\left\{\alpha\in L^{2}\Big{(}\mathbb{S}^{3};\bigwedge\nolimits^{2}T^{*}\mathbb{S}^{3}\Big{)}\colon d\alpha=0\,,Q(\alpha)>0\right\}\,,$ (4.2)
@@
$\mathcal{E}_{\rho}(\alpha):=\frac{2}{Q(\alpha)^{\frac{1}{2}}}\int_{\mathbb{S}^{3}}|\alpha|+\frac{\rho^{-2}}{Q(\alpha)}\int_{\mathbb{S}^{3}}|\alpha|^{2}\,.$ (4.3)
@@
$\mathcal{FS}_{\rho}(u)\geq\mathcal{E}_{\rho}(u^{*}\omega)\,,$ (4.4)
@@
An important property of the relaxed energy $\mathcal{E}_{\rho}$ is its local $L^{2}$-coercivity at $E_{0,1}^{+}$, as asserted in the following proposition, which is similar to [17, Lemma 3.2] but is valid in an $L^{2}$-neighborhood of $E_{0,1}^{+}$, instead of a $C^{0}$-neighborhood as in [17].
@@
For every $\rho_{0}\in(0,\sqrt{2})$ there exist $\varepsilon_{0}:=\varepsilon_{0}(\rho_{0})>0$ and $c:=c(\rho_{0})>0$, such that the following statement holds. If $\rho\in(0,\rho_{0}]$ and $\alpha\in\mathcal{A}_{\mathcal{FS}}$ is such that $Q(\alpha)=1$ and
$\|\alpha-\alpha^{+}_{0,1}\|_{L^{2}(\mathbb{S}^{3})}=\textup{dist}_{L^{2}}(\alpha,E^{+}_{0,1})\leq\varepsilon_{0}\,,$ (4.7)
where $\alpha^{+}_{0,1}\in E^{+}_{0,1}$, then
$\mathcal{E}_{\rho}(\alpha)-\mathcal{E}_{\rho}(\alpha^{+}_{0,1})\geq c\big{\|}\alpha-\alpha^{+}_{0,1}\big{\|}_{L^{2}(\mathbb{S}^{3})}^{2}\,.$ (4.8)
@@
The normalization condition for the Hopf invariant in (4.5), together with the facts that $\alpha^{+}_{0,1}$ is a constant coefficient 2-form (see (4.6) and Remark 2.3), that

### v2

@@ 2.3. The Hopf map and horizontal conformality
In this subsection we gather some well-known facts about the Hopf map. We first recall the following definition, which the reader can find more generally in [baird2003harmonic, Definition 2.4.2 and Lemma 2.4.4]:
@@
$du(x)\circ du(x)^{t}=\frac{1}{2}|du(x)|^{2}\,\mathrm{Id}_{T_{u(x)}{\mathbb{S}}^{2}}\quad\text{for }\mathcal{H}^{3}\text{-a.e. }x\in{\mathbb{S}}^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}3\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}\,$, (2.13)
@@
The following lemma gives a simple but very useful pointwise inequality related to weak horizontal conformality. Its proof is elementary and can be found for instance in [isobe2008, Proof of Lemma 2.1].
@@
We remark that this type of estimate also plays an important role in the minimality arguments in [CoronGulliver].
@@
One can choose an orthonormal frame $\{\tau_{1},\tau_{2},\tau_{3}\}$, defined at a chart point $(t,e^{i\varphi_{1}},e^{i\varphi_{2}})\in{\mathbb{S}}^{3}$ as
@@ 3. The Hopf invariant for maps of finite Faddeev–Skyrme energy
In this section we prove some results concerning the definition and integrality of the Hopf invariant for maps with finite Faddeev–Skyrme energy, cf. (1.2). The integrality property of the Hopf invariant in this class was correctly claimed in [lin2004existence, Section 3] but, as explained in [auckly2010pontrjagin, §2.1.4], there is a gap in the argument. In [auckly2010pontrjagin, Corollary 5.2] a complete proof of integrality was given, but the proof therein is fairly long and not self-contained, as it relies on the earlier results by the authors in [Auckly2005]. Here we present a new, more concise proof of this integrality which, although also not self-contained, we believe to be fairly transparent, and which contains ingredients which will also be important in the sequel.
@@
and in view of the fact that $H^{2}_{\rm{dR}}({\mathbb{S}}^{3})=0$ there exists $\beta\in\Omega^{1}({\mathbb{S}}^{3})$ such that $u^{*}\omega_{{\mathbb{S}}^{2}}=d\beta$. The Hopf invariant or Hopf charge of $u$, see [whitehead1947expression], is then given by the integral formula (1.3), and in particular $Q(u)\in\mathbb{Z}$.
@@
$\mathcal{A}_{3}({\mathbb{S}}^{3}):=\big\{v\in W^{1,2}({\mathbb{S}}^{3};{\mathbb{S}}^{3})\colon\textup{cof}(dv)\in L^{3/2}({\mathbb{S}}^{3})\big\}\,$,
where $\textup{cof}(dv)$ is identified with the pull-back map $v^{*}\colon\Omega^{2}({\mathbb{S}}^{3})\to\Omega^{2}({\mathbb{S}}^{3})$. These function spaces were introduced in the context of nonlinear elasticity theory in [Ball1981a] and appear naturally in our problem, since we have the following generalization of the results in [Universite1992, Muller1994, Sverak1988], due to [Hamburger1999, Theorems 1.3 and 4.4]:
@@
[…]
In addition to Proposition 3.2, the following result due to [Bethuel1991] is the key ingredient in our proof of Theorem 3.3.
@@
By [Bethuel1991, Theorem 1], a map $u\colon{\mathbb{S}}^{3}\to{\mathbb{S}}^{2}$ is $W^{1,2}$-strongly approximable by smooth maps if and only if it pulls-back $2$-forms on ${\mathbb{S}}^{2}$ to (distributionally) closed 2-forms on ${\mathbb{S}}^{3}$. Since any form in $\Omega^{2}(\mathbb{S}^{2})$ can be represented as $f\omega_{\mathbb{S}^{2}}$ for some $f\in C^{\infty}({\mathbb{S}}^{2})$, it is then necessary and sufficient to check closedness of $u^{*}\omega_{\mathbb{S}^{2}}$. ∎
@@
This is detailed in [auckly2010pontrjagin, Lemma 3.2]. Fix $\{j,k,\ell\}\subset\{1,2,3\}$. Thanks to classical approximation by (localization and) mollification, we see that $u^{k},u^{\ell}\in W^{1,2}({\mathbb{S}}^{3};\mathbb{R})$ satisfy
@@
hence the 2-form $\alpha^{k\ell}\colon=du^{k}\wedge du^{\ell}$ is distributionally closed, and using again approximation by a mollification of $u^{j}\in W^{1,2}({\mathbb{S}}^{3};\mathbb{R})$ and of the closed form $\alpha^{k\ell}\in L^{2}({\mathbb{S}}^{3};\bigwedge^{2}T^{*}{\mathbb{S}}^{3})$, we infer that
@@
$d(u^{*}\omega_{\mathbb{S}^{2}})=3\,du^{1}\wedge du^{2}\wedge du^{3}\in L^{1}\Big({\mathbb{S}}^{3};\bigwedge\nolimits^{3}T^{*}{\mathbb{S}}^{3}\Big)\,$, (3.4)
@@
Since $u(x)\in\mathbb{S}^{2}$ for $\mathcal{H}^{3}$-a.e. $x\in{\mathbb{S}}^{3}$, which implies
$u^{1}du^{1}+u^{2}du^{2}+u^{3}du^{3}=0\ \quad\text{for }\mathcal{H}^{3}\text{-a.e. }x\in{\mathbb{S}}^{3}\,$,
the vectors $(du^{i}(x))_{i=1}^{3}$ must be linearly dependent, hence the wedge product in the right hand side of (3.4) vanishes. ∎
The next lemma, due to [Hardt2003], gives us a gauge for the lifts of $u$.
@@
$h^{*}\omega_{{\mathbb{S}}^{2}}=2d\theta\,.$
@@
The proof in the case of a smooth map $u$ can be found in [Hardt2003, Lemma 2.1], and otherwise the conclusion of (3.6) can be repeated analogously, since the calculations therein can be carried over pointwise $\mathcal{H}^{3}$-a.e. on ${\mathbb{S}}^{3}$. The coefficients in (3.6) are actually different in [Hardt2003] due to a different choice of normalization for the metric on $\mathbb{S}^{2}$ (which also changes the normalization constant in the integral formula (1.3) for the Hopf degree). For the readers’ convenience we reproduce the proof of (3.6) with the metric used here. In the orthonormal frame $\{\tau_{1},\tau_{2},\tau_{3}\}$ on $T{\mathbb{S}}^{3}$, the $1$-form $\theta$ of (2.15) corresponds to the scalar product against $\tau_{1}$, cf. (2.19). Thus, the condition $\beta=2\hat{u}^{*}\theta$ and (2.18), (2.19), imply
@@
[…]
@@
since $\frac{5}{4}<\frac{3}{2}=:1^{*}$. We note that the borderline $L^{3/2}$-estimate is actually also true, see e.g. [LanzaniStein2005] or [BourgainBrezis2007, Corollary 12], at least in $\mathbb{R}^{3}$ instead of $\mathbb{S}^{3}$. Invoking Lemma 3.6, we now fix lifts $(\hat{u}_{j})_{j\in\mathbb{N}}\subset C^{\infty}({\mathbb{S}}^{3};{\mathbb{S}}^{3})$ satisfying
@@
cf. (3.5). In view of (2.14) and (3.9)-(3.11), by passing to a non-relabelled subsequence, we obtain $\beta\in L^{5/4}\big({\mathbb{S}}^{3};\bigwedge\nolimits^{1}T^{*}{\mathbb{S}}^{3}\big)$ so that
@@
Indeed, in view of (2.15), the forms $\hat{u}_{j}^{*}\theta$ can be written in coordinates as second-order polynomials of the form $\hat{u}_{j}\bullet d\hat{u}_{j}$, where the latter notation indicates that each monomial in the expression is of first order separately in $\hat{u}_{j}$ and $d\hat{u}_{j}$. In particular, since $\hat{u}_{j}\rightarrow\hat{u}$ strongly in $L^{p}({\mathbb{S}}^{3};{\mathbb{S}}^{2})$ for every $1\leq p<+\infty$ and $d\hat{u}_{j}\rightharpoonup d\hat{u}$ weakly in $L^{5/4}({\mathbb{S}}^{3};\bigwedge\nolimits^{1}(T^{*}{\mathbb{S}}^{3}\otimes\mathbb{R}^{3}))$, this particular product structure of weakly convergent objects and strongly convergent ones justifies (3.16).
@@
where $\{f_{1},f_{2}\}$ is the orthonormal frame on $T{\mathbb{S}}^{2}$ defined after (2.20). Thus, using also (2.4), we have
@@
[…]
@@
[…]
@@
There is a serious obstruction to having a more direct proof of Theorem 3.3, namely that it is not known whether one can approximate in the Faddeev–Skyrme energy a general map in $\mathcal{U}_{\mathcal{FS}}$ by smooth maps. This is not known even if one ignores the image constraint on maps in $\mathcal{U}_{\mathcal{FS}}$, and we refer the reader to [Hajasz2002] for some partial results in the unconstrained case. We emphasize that from the above proof one cannot infer that the Hopf invariant of $u$ is obtained as the limit of the Hopf invariants of $(u_{j})_{j\in\mathbb{N}}$ as in (3.9), since the latter sequence converges to $u$ only in too weak a sense for this purpose (and similarly for the liftings).
@@
Since the map $H^{2}_{\mathrm{dR}}({\mathbb{S}}^{2})\ni[\omega]\mapsto\int_{{\mathbb{S}}^{2}}\omega\in\mathbb{R}$ is a linear isomorphism, our assumption (3.18) in particular implies that
$\omega_{1}]=[\omega_{2}]\in H_{\mathrm{dR}}^{2}({\mathbb{S}}^{2})\,$,
@@
$=\int_{{\mathbb{S}}^{3}}\beta_{2}\wedge d\beta_{2}+2\int_{{\mathbb{S}}^{3}}d\beta_{2}\wedge v^{*}\eta+\int_{{\mathbb{S}}^{3}}v^{*}\eta\wedge v^{*}d\eta$
@@
$\int_{{\mathbb{S}}^{2}}\omega_{1}=\int_{{\mathbb{S}}^{2}}\omega_{2}\,.$
@@ 4. Relaxed form of the perturbed Faddeev–Skyrme energy
In order to prove Theorem 1.1, and following the approach in [isobe2008, riviere1998, riviere1998remark], we consider a relaxation of $\mathcal{FS}_{\rho}$ to the space of closed differential forms. Let us first define, for a closed 2-form $\alpha$ on ${\mathbb{S}}^{3}$,
@@
By Stokes’ Theorem, this quantity, referred to as the Hopf invariant of the form $\alpha$, is well-defined, i.e., $Q$ depends on $\beta$ only through $d\beta$, see also the corresponding discussion in Remark 3.1 for the Hopf invariant of a map $u$. Introducing the space of admissible forms
$\mathcal{A}_{\mathcal{FS}}:=\left\{\alpha\in L^{2}\Big(\mathbb{S}^{3};\bigwedge\nolimits^{2}T^{*}\mathbb{S}^{3}\Big)\colon d\alpha=0\,,Q(\alpha)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=1\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\right\}\,$,
@@
[…]
@@
$\mathcal{FS}_{\rho}(u)\geq\mathcal{E}_{\rho}(u^{*}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{{\mathbb{S}}^{2}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})\,$,
@@
The next proposition for the relaxed energy $\mathcal{E}_{\rho}$ is the main ingredient in the proof of Theorem 1.1. Compared with [isobe2008, Lemma 3.2], Proposition 4.2 asserts global minimality of $E_{0,1}^{+}$, while the latter only gives minimality in a $C^{0}$-neighborhood of $E_{0,1}^{+}$.
@@
For every $\rho\in(0,1]$ and $\alpha\in\mathcal{A}_{\mathcal{FS}}$, if
$\alpha^{+}_{0,1}\in\underset{\widetilde{\alpha}\in E_{0,1}^{+}}{\mathrm{argmin}}\,\|\alpha-\widetilde{\alpha}\|_{L^{2}({\mathbb{S}}^{3})}\,$, (4.5)
then
[…]
with equality if and only if $\alpha\in E_{0,1}^{+}$.
@@
The normalization condition for the Hopf invariant in (4.3), together with the facts that $\alpha^{+}_{0,1}$ is a constant coefficient 2-form, cf. (4.4) and Remark 2.3, that